# Filtering for Creativity: Adaptive Prompting for Multilingual Riddle Generation in LLMs

### v1

@@ A.1 English
As shown in Table 8, Row 1, the fine‑tuned riddle reimagines the original with more abstract and layered associations. Rather than relying on negated literalism, it introduces concepts like memory and time using metaphorical compression and cross-sensory cues. This approach reflects principles of conceptual integration theory, where blending disparate domains enhances figurative depth Fauconnier & Turner (2002). In contrast, the real-world version is more direct, using structural opposition to achieve its effect Gentner (1983).Row 2 presents another clear shift in stylistic strategy. The real-world riddle uses static reversal—a common riddle trope—while the fine-tuned variant introduces paradox and disappearance as metaphors for guidance. This relies on spatial embodiment, a known technique in metaphor production Lakoff & Johnson (1980), and adds ambiguity through indirectness. Such compositional layering suggests that AOF prompts can foster more interpretively open and stylistically creative riddles than conventional examples.
@@ A.2 Japanese
The riddles in AOF are guided towards direct metaphors with complex, creative, and unique word choice and sentence structure, while having creative answers like memory and beehive in Table 12 Teng & Xu (2023). These generations surpass past riddle generations flaws like lack of originality in sentence structure, just changing the pronouns or verbs to make it more creative, and etc. These riddles contrast with traditional Japanese riddles which rely on phonetic ambiguity and cultural nuance like in Table 12 where the first row features how phonetically similar words feature different meanings and the riddle in the second row yields different ways of reading through phonetically similar readingsAn (2023).
@@ A.3 Chinese
Fine-tuned AOF riddles in Chinese often leverage character structure through radical-based puns and vivid imagery. For instance, the coral riddle in Table 15 blends “sea” imagery with radical hints (海底藏森林…) to guide the solver—a strategy supported by prior work on character-pun alignments in riddle composition Tan et al. (2016). By contrast, traditional 灯谜 (e.g., “口袋里有个圆…” for “月亮”) rely on simple perceptual clues and tonal balance Wei & Lee (2021). This comparison suggests that our approach enhances cultural depth by embedding multi‑layered orthographic play into poetic metaphors while preserving reader accessibility.
A.4 Arabic
@@
(Figure 4, Row 5) AOF stands out for its fresh language and metaphorical clarity. One riddle—”Something that’s full when it eats, and thirsty when it drinks”—relies on a simple yet clever contradiction that invites reflection. It draws on the tradition of using everyday logic to confuse and amuse, evoking the style of oral riddles that play with basic physical experiences. The second riddle—”I light up the night and disappear by day, visible yet unseen… What am I?”—is more poetic, using contrast and imagery to express something elusive and symbolic. It captures the feel of classical Arabic alghāz not through root-based punning but through layered metaphor and rhythm. Together, these examples show how AOF preserves the spirit of traditional riddling through modern, metaphor-rich language Antar (2023); Bhatt & Kuka (2025); Liu et al. (2022).
A.5 French
@@
Fine-tuned AOF riddles in French lean into unexpected domain shifts and internal echo. The AOF example repurposes the concept of a “typo” as a buzzing bee, combining internal rhyme (“jardin/des mots”, “bourdonnant/lettres”) and metaphorical layering, driving semantic playfulness and rhythmic balance (Table 19, Row 1). Internal rhyme notably enhances poetic cohesion and cognitive engagement Encyclopædia Britannica (2025). In contrast, canonical French énigmes tend toward binary negation and elemental imagery (Table 19, Row 2). For instance, “Je vole sans ailes, je pleure sans yeux…” relies on simple antithesis without cross-domain metaphorical transfer. The AOF variant’s richer conceptual mapping aligns with findings that cross-domain metaphor and internal structure boost interpretability and novelty in poetic forms Lakoff & Johnson (1999); Encyclopædia Britannica (2025).
Appendix B Appendix: Additional Results Tables
B.1 Average Token Length Across Pretrained Models

### v3

@@ A.1 English
Fine-tuning GPT-4o with AOF notably improves semantic richness and lexical creativity (RiddleScore: 0.586; Table 4). AOF achieves superior lexical diversity (Distinct-2: 0.893) and minimal structural repetition (Self-BLEU: 0.260) compared to few-shot and adversarial baselines (Table 9), validating RiddleScore’s effectiveness as a comprehensive evaluation measure (Zhang et al., 2020b; Sellam et al., 2020b). Qualitatively, riddles such as those in Table 13 illustrate innovative metaphor usage and coherent ambiguity, consistent with cognitive theories on figurative language and memorability (Lakoff and Johnson, 1980; Koestler, 1964; Fauconnier and Turner, 2002). For instance, Row 1 deploys cues like "mirror yours" and "echo thoughts" to encode identity and perception into abstract form, while Row 2 evokes silence as an interstitial force through metaphors, aligning with conceptual blending theory (Fauconnier and Turner, 2002).
@@ A.2 Japanese
Fine-tuning GPT-4o with AOF significantly enhances morphosyntactic fluency and metaphor–answer cohesion in Japanese–English riddle generation (RiddleScore: 0.475; Table 4). Compared to other prompting methods, AOF produces riddles with the lowest structural redundancy (Self-BLEU: 0.177) and highest lexical diversity (Distinct-2: 0.915), indicating stronger semantic control and reduced overfitting to prior examples (Table 15). These gains are reflected in AOF’s leading RiddleScore, which surpasses Zero-Shot (0.300), Few-Shot (0.334), CoT (0.307), and Adversarial (0.331) settings. Qualitatively, the generated riddles exhibit hallmarks of Japanese poetic reasoning—syntactic compression, metaphorical layering, and rhythmical closure—without resorting to direct translation or formulaic repetition(Kawamura, 2016). For instance,「屋根にはいるのに、家にいないものは何?」(“What enters the roof but never the house?”) leverages spatial contradiction in a culturally familiar frame, while maintaining logical symmetry across both languages(Heine and Kuteva, 2007). This fidelity to both Japanese linguistic nuance and cross-lingual metaphor construction is characteristic of AOF’s superiority, suggesting greater alignment with human intuitions of creativity, fluency, and interpretability.
@@ A.3 Chinese
Fine-tuning enhances metaphorical control and orthographic awareness in Chinese riddles. AOF outputs consistently avoid overused oppositional templates like “我有…却…,” favoring layered metaphors, radical-based hints, and prosodic fluency. Compared to Zero-Shot and Few-Shot baselines, AOF achieves lower Self-BLEU (0.163 vs. 0.315 / 0.349) and higher Distinct-2 (0.934 vs. 0.831 / 0.787), validating RiddleScore as a composite indicator of structural novelty (0.728; Table 4(Zhang et al., 2020b; Sellam et al., 2020b). In Table 20, Row 1 evokes lunar imagery with rhythmic balance, updating a classical riddle (“口袋里有个圆…”) through spatial metaphor and contrast (Sun, 2006; Wei and Lee, 2021). Row 2 exemplifies orthographic metaphor: “信” is revealed through poetic compression (“千言万语藏心怀”), echoing traditional pun-encoding in radical-based 灯谜 (Tan et al., 2016; Li, 2008). Row 3 (蝴蝶) combines temporal framing and sensory motion (“彩衣…花丛…无踪”) to support multi-modal reasoning, in line with conceptual blending theory (Fauconnier and Turner, 2002; Lakoff and Johnson, 1980). These results suggest that AOF produces culturally grounded riddles with high interpretability and lexical range.
A.4 French
Fine‑tuning GPT‑4o with AOF yields French riddles that combine varied grammatical forms, fresh metaphors, and cultural resonance. The model moves beyond standard “Qu’est‑ce qui…” stems and elemental tropes to embrace declarative statements, poetic ellipses, and even modern imagery. Although Zero-Shot achieves a higher RiddleScore (0.468 vs. 0.352), AOF excels in lexical diversity (Distinct-2 = 0.856) and maintains moderate repetition (Self-BLEU = 0.273), suggesting greater creative variance in form and framing (Zhang et al., 2021; Binsted, 1996). AOF riddles (Table 24) avoid clichés like “ombre” or “écho” and instead draw on subtle metaphor and rhythm. For instance, Row 1 uses cyclical phrasing to express the return of day (jour), while Row 2 reframes a broom through trailing ellipsis and implied motion. These constructions echo prior findings on metaphor-induced novelty and poetic ambiguity (Lakoff and Johnson, 1980; Koestler, 1964), even when metric scores undervalue such stylistic range.
A.5 Arabic
Fine-tuning GPT-4o with AOF improves semantic richness and metaphorical ingenuity in Arabic–English bilingual riddles (RiddleScore: 0.586; Table 4). Compared to Few-Shot (0.364), Zero-Shot (0.315), Chain-of-Thought (0.313), and Adversarial (0.341), AOF achieves higher lexical variety (Distinct-2: 0.893) and lower repetition (Self-BLEU: 0.260), showing its balance between novelty and coherence (Table 9). These results confirm RiddleScore’s effectiveness for evaluating creativity and linguistic depth (Zhang et al., 2020c; Sellam et al., 2020b). Qualitatively, AOF riddles reflect traditional Arabic poetic traits—metaphorical layering, conceptual blending, and cultural framing—without relying on literal translation. For example, Figure 6, Row 1 uses sound as a metaphor for something intangible yet present—"I exist in the air, yet I do not fly"—echoing classical rhetoric. Row 2 likens strong wind to a guest who "passes nearby homes but is never welcome inside". These examples illustrate nuanced cultural imagery and poetic reasoning, consistent with the richness of Arabic literary tradition (Al-Khatib, 1988). AOF thus enhances both creativity and interpretability in bilingual Arabic riddles.
Appendix B Appendix: AOF Pretrained language Evaluations
B.1 English
GPT-4o
achieves moderate repetition (Self-BLEU: 0.413) and high lexical diversity (Distinct-2: 0.852), balancing structural cohesion with surface novelty. These characteristics correspond with its AOF RiddleScore of 0.373, indicating that while GPT-4o avoids excessive repetition, its metaphorical expressiveness remains moderate. Compared to LLaMA 3.1 (0.471 / 0.727, RiddleScore: 0.352) and DeepSeek R1 (0.339 / 0.845, RiddleScore: 0.400), GPT-4o represents a middle ground: less phrasally diverse than R1, but more structurally consistent than LLaMA. The riddle in Row 1 of Table 10 reflects these tendencies, blending contrastive metaphor with cohesive syntax. This supports prior findings that figurative ambiguity coupled with syntactic regularity enhances interpretability (Lakoff and Johnson, 1980; Shutova, 2013).
LLaMA 3.1
displays the strongest phrasal variation (Distinct-2: 0.727), but with moderately higher repetition (Self-BLEU: 0.471) and a slightly lower AOF RiddleScore of 0.352. These metrics suggest that while LLaMA 3.1 explores more varied lexical forms, it occasionally overuses structural templates. The riddle in Row 2 of Table 10 shows rhythmic symmetry and layered metaphor, reinforcing theories linking riddle memorability to structured cadence and salience (Koestler, 1964). The AOF prompt appears to mitigate lexical rigidity by encouraging recomposition within constrained semantic bounds (Fauconnier and Turner, 2002).
DeepSeek R1
demonstrates the lowest repetition (Self-BLEU: 0.339), highest lexical diversity (Distinct-2: 0.845), and the top AOF RiddleScore at 0.400, indicating superior expressive range and originality. The riddle in Row 3 exemplifies conceptual inversion, pairing abstract imagery with narrative misdirection—a hallmark of classic riddle mechanics (Koestler, 1964). While extreme novelty sometimes threatens fluency (Zhang et al., 2021), R1’s outputs remain syntactically intact, suggesting that AOF balances expressiveness with readability (Xu et al., 2018). This balance likely contributes to R1’s higher perceived riddle quality as measured by RiddleScore.
B.2 Japanese
GPT-4o
While GPT-4o’s performance on metrics like self-BLEU and distinct-n using the AOF prompt falls around the average compared to standard baselines, it excels notably in RiddleScore, achieving a score of 0.475. This substantial increase over traditional methods (Few-Shot: 0.334, Zero-Shot: 0.300, Chain-of-Thought: 0.307, Adversarial: 0.331) reflects the model’s ability to generate riddles with greater novelty, fluency, diversity, and semantic coherence ((Yao et al., 2025), (Schmidtová and Wu, 2024)). AOF specifically addresses traditional prompting flaws such as the "I"-centered imagery prevalent in chain-of-thought prompts and the example-specific overfitting observed in few-shot prompts, thereby substantially enhancing multilingual riddle quality. For instance, the riddle example in Table 14 features a distinctive structure—a concise opening followed by a more elaborate second sentence—which enhances reader engagement and contributes to its high RiddleScore.
LLaMa3.1
Although LLaMa3.1 does not demonstrate significant improvement in automated metrics like self-BLEU and distinct-n under the AOF framework, its RiddleScore of 0.475 significantly surpasses traditional baselines (Few-Shot: 0.334, Zero-Shot: 0.300, Chain-of-Thought: 0.307, Adversarial: 0.331). This highlights AOF’s effectiveness in enhancing multilingual riddle generation beyond conventional evaluation metrics by addressing issues such as egocentric phrasing and repetition. Notably, the riddle presented in Table 14 cleverly employs the homophone 「つる」, invoking both decorative twine and the crane (鶴)—elements deeply embedded in Japanese cultural symbolism and Shinto rituals like しめ縄 (shimenawa) (An, 2023). This cultural and linguistic depth significantly contributes to its superior RiddleScore.
DeepSeek R1
DeepSeek R1, while only achieving median results on surface-level metrics such as self-BLEU and distinct-n, shows marked improvement with a RiddleScore of 0.475 compared to lower scores from standard methods (Few-Shot: 0.334, Zero-Shot: 0.300, Chain-of-Thought: 0.307, Adversarial: 0.331). The RiddleScore clearly underscores the efficacy of the AOF prompting strategy in overcoming baseline shortcomings like excessive first-person imagery and rigid replication patterns, promoting originality, fluency, and semantic coherence. An illustrative example from Table 14 artfully misleads readers by metaphorically describing a fish’s mouth as a "quiet tree" where birds sing, skillfully blending surreal imagery with natural elements(DiStefano and Patterson, 2024). This innovative poetic device significantly enhances its overall RiddleScore.
B.3 Arabic
GPT-4o
GPT-4o shows moderate repetition (Self-BLEU: 0.497) and good lexical variety (Distinct-2: 0.780) with Adaptive Originality Filtering (AOF), clearly performing better than common methods like few-shot, zero-shot, chain-of-thought, and adversarial prompts.With an AOF RiddleScore of 0.373, GPT-4o demonstrates notable improvement over chain-of-thought (0.304) and adversarial methods (0.296).Unlike chain-of-thought prompts, which tend to produce straightforward, predictable metaphors, AOF helps GPT-4o create riddles with imaginative and abstract images—such as something that’s present but unseen—as illustrated in (Figure 6, Row 1). This approach fits naturally with traditional Arabic riddles, known for their symbolic and reflective style (Al-Khatib, 1988).
LLaMA 3.1
LLaMA 3.1 strikes an effective balance between repetition (Self-BLEU: 0.374) and creativity (Distinct-2: 0.927) through AOF, resulting in a RiddleScore of 0.378. This addresses issues often found in chain-of-thought (0.303) and adversarial prompts (0.292), which frequently yield predictable or overly vague outputs. Its riddles are relatable and culturally resonant, using clear metaphors drawn from everyday life, like "a strong wind" that can’t enter a house, as shown in (Figure 6, Row 2). This connects directly to familiar poetic traditions in Arabic, avoiding common pitfalls like repetitive phrasing or loss of meaning (Al-Jahiz, 869).
DeepSeek R1
DeepSeek R1, while somewhat repetitive (Self-BLEU: 0.585), achieves notable depth in metaphorical expression (Distinct-2: 0.583) under AOF, resulting in the highest RiddleScore of 0.400 among the three models. This method effectively tackles problems seen in zero-shot (0.400), few-shot (0.341), chain-of-thought (0.304), and adversarial prompting (0.305), such as repetitive or simplistic metaphors. For example, DeepSeek R1 creatively portrays a rooftop as an eye "fed by the city," as seen in (Figure 6, Row 3), mixing urban imagery with striking visual symbolism. This clever blending of abstract ideas and real-world images strongly aligns with Arabic poetry, known for its layers of meaning and subtle metaphors (Al-Marzouki, 2012). By encouraging culturally rich riddles, AOF clearly boosts the originality and depth of DeepSeek R1’s outputs compared to simpler prompting strategies (Xu et al., 2018).
B.4 French
GPT-4o
GPT-4o’s pretrained riddles are grammatically fluent and consistently answerable, but often exhibit translated literalism rather than native poetic expressivity. For instance, its output in Row 1 of Table 21 invokes elemental imagery typical of English-origin riddles, but lacks stylistic markers common in French verse, such as enjambment or internal rhyme (Delisle, 1999). These tendencies yield a Self-BLEU of 0.413 and a high Distinct-2 of 0.852, suggesting strong surface diversity but moderate structural reuse. This balance corresponds to an AOF RiddleScore of 0.373, reflecting a safe, comprehensible style with limited cultural specificity or rhythmic nuance (Chan, 1996).
DeepSeek R1
DeepSeek R1 offers concise and semantically transparent riddles, often echoing patterns from elementary French folklore. As seen in Row 2, its outputs favor concrete dualities (“bed but never sleep”) common in children’s riddles (Meulemans, 2005), yielding low Self-BLEU (0.339) and high Distinct-2 (0.845). These surface metrics align with an AOF RiddleScore of 0.354, indicating moderate creativity tempered by formulaic structure. While effective, R1’s riddles seldom explore prosodic depth or figurative abstraction (Leman, 2013), limiting their stylistic innovation despite syntactic precision.
LLaMA 3.1
LLaMA 3.1 demonstrates the widest stylistic bandwidth among pretrained models. Its Row 3 output juxtaposes dance and laughter through internal echo, while Row 4 ventures into digital metaphor with a riddle about a cursor. These examples reflect the model’s capacity for modernized symbolic extension, albeit inconsistently. With a Self-BLEU of 0.471, Distinct-2 of 0.727, and RiddleScore of 0.352, LLaMA balances lexical innovation with occasional overreach. These fluctuations suggest strong creative potential but uneven cohesion, echoing prior observations on metaphor blending and linguistic recombination (Veale, 2011; Binsted, 1996).
B.5 Chinese
GPT-4o
GPT-4o’s pretrained Chinese riddles are grammatically correct and logically coherent, but often translate English metaphors without adapting to the script-specific strategies typical of traditional 灯谜. As shown in Row 1 of Table 17, the imagery is literal and binary, missing multi-layered allusions like radical-based clues or idiomatic rhythm (Chan, 1996; Sun, 2006). With a Self-BLEU of 0.280, Distinct-2 of 0.869, and an AOF RiddleScore of 0.434, the model achieves surface novelty without fully leveraging character-level poetic mechanisms. This suggests competent fluency but limited cultural depth.
DeepSeek R1
DeepSeek R1 produces elegant, fluent couplets with classical poetic symmetry, as seen in Row 2. While rhythm and antithesis are preserved, metaphors remain literal—favoring structural form over layered meanings. This is reflected in a Self-BLEU of 0.433, Distinct-2 of 0.674, and an AOF RiddleScore of 0.453, the highest among the three models. The results indicate that while R1 may lack idiomatic richness, it effectively balances structural clarity and lexical diversity, offering consistently coherent outputs with stylistic restraint (Xu et al., 2018).
LLaMA 3.1
LLaMA 3.1 exhibits the richest cultural range in pretrained generation. Row 4 blends visual and semantic metaphor reminiscent of folk riddles, and Row 5 demonstrates radical-based structure. Its Distinct-2 of 0.776 and Self-BLEU of 0.428 align with an AOF RiddleScore of 0.330, revealing moderate creativity yet lower overall cohesion. Although stylistically ambitious, LLaMA occasionally struggles with logic or phrasing. Still, its outputs reflect deeper integration with Chinese morphological conventions than its counterparts (Li, 2008; Fauconnier and Turner, 2002).
Appendix C Appendix: Fine-Tuned AOF Riddle Comparison to Real World
C.1 English
As shown in Table 12, Row 1, the fine‑tuned riddle reimagines the original with more abstract and layered associations. Rather than relying on negated literalism, it introduces concepts like memory and time using metaphorical compression and cross-sensory cues. This approach reflects principles of conceptual integration theory, where blending disparate domains enhances figurative depth (Fauconnier and Turner, 2002). In contrast, the real-world version is more direct, using structural opposition to achieve its effect (Gentner, 1983).Row 2 presents another clear shift in stylistic strategy. The real-world riddle uses static reversal—a common riddle trope—while the fine-tuned variant introduces paradox and disappearance as metaphors for guidance. This relies on spatial embodiment, a known technique in metaphor production (Lakoff and Johnson, 1980).
C.2 Japanese
The riddles in AOF are guided towards direct metaphors with complex, creative, and unique word choice and sentence structure, while having creative answers like memory and beehive in Table 16 (Teng and Xu, 2023). These generations surpass past riddle generations flaws like lack of originality in sentence structure, just changing the pronouns or verbs to make it more creative, and etc. These riddles contrast with traditional Japanese riddles which rely on phonetic ambiguity and cultural nuance like in Table 16 where the first row features how phonetically similar words feature different meanings and the riddle in the second row yields different ways of reading through phonetically similar readings(An, 2023).
C.3 Chinese
@@
Fine-tuned AOF riddles in Chinese often leverage character structure through radical-based puns and vivid imagery. For instance, the coral riddle in Table 19 blends “sea” imagery with radical hints (海底藏森林…) to guide the solver—a strategy supported by prior work on character-pun alignments in riddle composition (Tan et al., 2016). By contrast, traditional 灯谜 (e.g., “口袋里有个圆…” for “月亮”) rely on simple perceptual clues and tonal balance (Wei and Lee, 2021). This comparison suggests that our approach enhances cultural depth by embedding multi‑layered orthographic play into poetic metaphors while preserving reader accessibility.
C.4 Arabic
@@
(Figure 8, Row 5) AOF stands out for its fresh language and metaphorical clarity. One riddle—"Something that’s full when it eats, and thirsty when it drinks"—relies on a simple yet clever contradiction that invites reflection. It draws on the tradition of using everyday logic to confuse and amuse, evoking the style of oral riddles that play with basic physical experiences. The second riddle—"I light up the night and disappear by day, visible yet unseen… What am I?"—is more poetic, using contrast and imagery to express something elusive and symbolic. It captures the feel of classical Arabic alghāz not through root-based punning but through layered metaphor and rhythm. Together, these examples show how AOF preserves the spirit of traditional riddling through modern, metaphor-rich language (Antar, 2023; Bhatt and Kuka, 2025; Liu et al., 2022b).
C.5 French
@@
Fine-tuned AOF riddles in French lean into unexpected domain shifts and internal echo. The AOF example repurposes the concept of a “typo” as a buzzing bee, combining internal rhyme (“jardin/des mots”, “bourdonnant/lettres”) and metaphorical layering, driving semantic playfulness and rhythmic balance (Table 23, Row 1). Internal rhyme notably enhances poetic cohesion and cognitive engagement (Encyclopædia Britannica, 2025). In contrast, canonical French énigmes tend toward binary negation and elemental imagery (Table 23, Row 2). For instance, “Je vole sans ailes, je pleure sans yeux…” relies on simple antithesis without cross-domain metaphorical transfer. The AOF variant’s richer conceptual mapping aligns with findings that cross-domain metaphor and internal structure boost interpretability and novelty in poetic forms (Lakoff and Johnson, 1999; Encyclopædia Britannica, 2025).
Appendix D Appendix:Fine-Tuned vs. Pretrained Riddle Generation
282830303232343436363838404042424444464648485050525254545656585805510101515ArabicChineseEnglishFrenchJapaneseImprovement in RiddleScore (%)Improvement in Human Evaluation (%)ArabicChineseEnglishFrenchJapanese
Figure 4: Correlation between fine-tuning gains in RiddleScore and human evaluation scores across five languages. Each point represents one language; higher values correspond to more improvement compared to the pretrained model.
ArabicChineseEnglishFrenchJapanese50-500505010010057.157.148.348.343.443.433.733.729.529.533.5-33.551.3-51.340.4-40.443.2-43.263.4-63.418.818.810.610.613.513.513.313.331.331.3015.115.116.916.91.3-1.32.12.1Change from Pretrained (%)RiddleScore GainSelf-BLEU ReductionDistinct-2 GainHuman Eval Gain
Figure 5: Percentage changes in RiddleScore, Self-BLEU, Distinct-2, and human evaluation after fine-tuning. Positive bars show improvements; negative Self-BLEU values (in red) indicate desirable reductions in repetition.
We compare GPT-4o before and after fine-tuning across five prompting strategies. Quantitative metrics—token length, Self-BLEU, and Distinct-2—are complemented by qualitative analysis of metaphorical framing, structural variation, and bilingual phrasing. Representative pairs are shown in Appendices GJ.
D.1 English
Fine-tuning reduces token length by 28.2%, repetition by 40.4%, and increases lexical variety by 6.1%. RiddleScore improves by 43.4%, showing that reduced redundancy and more diverse phrasing lead to higher-quality riddles. Pretrained outputs often reflect familiar patterns like personification, while fine-tuned ones adopt more abstract and fluent structures (Table 11, Row 1). Few-shot fine-tuning increases metaphorical expression but also length, with a 44.9% gain in RiddleScore (Row 2). CoT prompts benefit most—token length drops by 37.6%, diversity rises by 13.6%, and RiddleScore jumps 48.5% (Row 3). AOF produces the most creative riddles with metaphors like “quietest word,” improving RiddleScore by 42.9% alongside strong gains in novelty and clarity. Adversarial fine-tuning increases abstraction while reducing repetition by 18.2%, improving lexical diversity by 9.4%, and boosting RiddleScore by 33.4% (Row 5) (Zhang et al., 2020b; Sellam et al., 2020b).
D.2 Japanese
Across all prompting methods, fine-tuning improves morphosyntactic fluency and metaphorical layering. In Zero-Shot (Table 15, Row 1), outputs drop by 15.6% in Self-BLEU and align better with Japanese poetic rhythm(Kojima et al., 2022). Few-shot prompts (Table 15, Row 2) benefit from clearer clause structure and cultural framing, resulting in a 28.6% increase in distinct-n. CoT outputs (Table 15, Row 3) shift from templated “I…” forms to more idiomatic bilingual logic, improving Self-BLEU by 27.5% and 27.4% shorter riddles on average. Adversarial riddles (Table 15, Row 4) gain fluency and metaphor variation while reducing structural awkwardness. Yet, across Zero-Shot, Few-Shot, and CoT prompting, RiddleScore remained largely unchanged when moving from the pretrained to the fine-tuned model, suggesting that improvements in fluency and metaphorical richness did not translate into deeper semantic cohesion(Resck et al., 2024). Notably, Adversarial prompting saw a 14.5% drop in RiddleScore, indicating that its gains in stylistic fluency and metaphor density may have come at the cost of the semantic originality and structural coherence captured by the metric(Resck et al., 2024). In contrast, AOF prompting (Table 15, Row 5) exhibited no such trade-off, achieving the largest qualitative gain: a 63.4% drop in Self-BLEU, 31.3% increase in Distinct-2, and a 29.5% improvement in RiddleScore, reflecting enhanced metaphor density and cultural cadence without sacrificing semantic quality.
D.3 Chinese
Fine-tuning improves both variety and clarity in riddle phrasing. In Zero-Shot (Table 18, Row 1), replacing rigid sentence frames lowers repetition by 6.0%, boosts lexical diversity by 3.0%, and improves RiddleScore by 6.5%. Few-shot fine-tuning (Row 2) preserves strong metaphor use while avoiding repeated idioms, improving RiddleScore by 14.7%, increasing diversity by 5.6%, and reducing repetition by 6.8%. CoT prompts (Row 3) yield shorter riddles with smoother structure, cutting token length by 26.4%, increasing Distinct-2 by 7.2%, and raising RiddleScore by 18.1%. Adversarial fine-tuning (Row 4) boosts rhythm and cohesion, increasing lexical variety by 9.3% and RiddleScore by 12.8%, despite a 10.2% rise in repetition. AOF (Row 5) produces the most abstract and fluent riddles, lowering repetition by 51.3%, raising diversity by 10.6%, shortening outputs by 25.1%, and improving RiddleScore by 48.3% (Zhang et al., 2020b; Sellam et al., 2020b).
D.4 Arabic
Fine-tuning significantly enhances lexical diversity, reduces redundancy, and improves riddle quality in Arabic. In Zero-Shot (Table 7, Row 1), fine-tuned riddles replace rigid "X without Y" structures with rhythmic phrasing, reducing repetition (Self-BLEU) by 33.5%, increasing lexical diversity (Distinct-2) by 18.8%, and enhancing RiddleScore by 10.5% (0.315 to 0.348). Few-shot prompts (Row 2) abandon repetitive frames for enjambment and root variation, reducing Self-BLEU by 5.6%, increasing Distinct-2 by 8.1%, and improving RiddleScore by 8.0% (0.364 to 0.393). Chain-of-Thought (CoT) riddles (Row 3) become concise and idiomatic, lowering redundancy by 21.0%, increasing lexical diversity by 2.8%, and improving RiddleScore by 3.6% (0.313 to 0.324). Adversarial prompting (Row 4) introduces triadic parallelism and poetic misdirection, substantially reducing repetition by 44.7%, boosting lexical variety by 13.0%, and raising RiddleScore by 15.2% (0.341 to 0.393). AOF (Row 5) maintains peak lexical diversity (18.8% increase), decreases redundancy by 33.5%, and achieves the highest RiddleScore improvement (57.1%; 0.373 to 0.586), aligning closely with traditional Arabic poetic conventions (Al-Khatib, 1988).
D.5 French
Fine-tuning reduces dependence on literal templates like “Qu’est-ce qui…” and improves vocabulary variety across all prompt styles. In Zero-Shot (Table 22, Row 1), riddles shift from repetitive phrases to richer idiomatic expressions, with a 43.2% drop in Self-BLEU and a 7.1% rise in Distinct-2. RiddleScore improves by 25.7%, reflecting increased originality. Few-shot prompts (Row 2) yield riddles that are 32.8% shorter, with a 20.7% reduction in repetition and a 9.7% boost in lexical diversity; RiddleScore climbs 26.0%. CoT (Row 3) strikes a strong balance: repetition drops by 26.6%, diversity improves by 13.3%, and RiddleScore rises by 32.5%. Adversarial prompting (Row 4) enhances clarity while preserving misdirection, yielding a 14.0% reduction in Self-BLEU, 7.6% gain in Distinct-2, and 24.1% improvement in RiddleScore. AOF (Row 5) performs best overall, cutting repetition by 42.4%, achieving peak diversity, and delivering a 33.7% boost in RiddleScore. These results suggest that reducing redundancy and using more expressive, domain-appropriate language leads to riddles that are more fluent and culturally aligned (Zhang et al., 2020b; Sellam et al., 2020b).
Appendix E Appendix: Additional Results Tables
E.1 Average Token Length Across Pretrained Models